OpenStudy (anonymous):

Choose the equation below whose axis of symmetry is x = 2. Answer y = x2 + 4x + 2 y = x2 - 4 y = x2 - 2 y = x2 - 4x + 2

OpenStudy (anonymous):

x2 - 4x + 2 cuzz f(x-4)= (x-4)^2 expand and (x-4)(x-4) = that :P

OpenStudy (anonymous):

to double check x^2 - 4*x + 2 plug it into google